OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? Calculate the volume of a sphere with a radius of 3 in. Express your answer in terms of π. V = 36π in.3 a. 36π in.3 b. 45π in.3 c. 63π in.3 d. 81π in.3

OpenStudy (anonymous):

V=4/3 pi r^3

OpenStudy (anonymous):

V=4/3 (3)^3 pi =4*9 pi=36 pi in^3

OpenStudy (anonymous):

you are right A is the answer

OpenStudy (anonymous):

\[4/3 \times \pi \times r^3\]

OpenStudy (anonymous):

r is 3 \[4/3 \times \pi \times 3^3\]

OpenStudy (anonymous):

\[3^3 = 3 \times 3 \times 3 = 9 \times 3 = 27\]

OpenStudy (anonymous):

so now \[4/3 \times \pi \times 27\]

OpenStudy (anonymous):

now \[4/3 \times 27\]

OpenStudy (anonymous):

so 3 times 9 is 27 so now \[4 \times 9 \]

OpenStudy (anonymous):

which is \[4 \times 9 = 36\]

OpenStudy (anonymous):

therefore out of the options given A is correct